1.) Consolidated Software​ doesn't currently pay any dividends but is expected to start doing so in 4 years. That​ is, Consolidated will go 3 more years without paying any dividends and then is expected to pay its first dividend​ (of $1.41 per​ share) in the fourth year. Once the company starts paying​ dividends, it's expected to continue to do so. The company is expected to have a dividend payout ratio of 41​% and to maintain a return on equity of 21​%. Based on the​ DVM, and given a required rate of return of 16​%, what is the maximum price you should be willing to pay for this stock​ today?

2.) Assume you obtain the following information about a certain​ company:

Total assets

​$51,000,000

Total equity

​$21,390,984

Net income

​$2,195,158

EPS

​$2.64 per share

Dividend payout ratio

39​%

Required return

7.4​%

Use the​ constant-growth DVM to place a value on this​ company's stock.

3.) AviBank Plastics generated an EPS of ​$2.89 over the last 12 months. The​ company's earnings are expected to grow by 16.5% next​ year, and because there will be no significant change in the number of shares​ outstanding, EPS should grow at about the same rate. You feel the stock should trade at a​ P/E of around 19 times earnings. Use the​ P/E approach to set a value on this stock.

please help! I have no idea what I'm doing
Answer #1

Solution :- 1

Dividend in year 4 = $1.41

Here now we find growth rate (g) = b*r = (1-0.41)*0.21 = 0.1239 = 12.39%

Ke = 16%

Now Price at year 4 = D(1+g)/(ke-g) = 1.41(1+0.1239)/(0.16-0.124) = 1.584/0.0361 = 43.897$

Now price at year 0 = P4/(1+g)4 = 43.897/(1+0.1239)4

= $27.51

Solution 3 :-

Current EPS= $2.89

Expected EPS= $2.89* (1+ 16.5%) = $3.36685

Expected P/E Ratio = 19 times

Hence Price of the stock= PE Ratio * EPS = 19*3.3668 = $63.97

As per HOMEWORKLIB POLICY there is need to solve one answer at once even I answer 2 of these so please ask the 2nd question as separate one